How many moles are in 987.0 grams of Ra(OH)2


1
Expert's answer
2022-05-25T02:34:04-0400

m(Ra(OH)2) = 987.0 g

M(Ra(OH)2) = 260.o g/mol

m = n*M

n = m/M = 987.0 g / 260.o g/mol = Ra(OH)2)

Answer:

n(Ra(OH)2) = Ra(OH)2)